OpenStudy (anonymous):

****************************Medals and fans**************************************************At the beginning of the week, Paloma opened a new box of 100 coffee creamers. Forty-seven coffee creamers were used during the week. Which equation could be used to find x, the number of coffee creamers left at the end of the week?

OpenStudy (anonymous):

47 - 100 = x 47x = 100 x + 47 = 100 x - 47 = 100

OpenStudy (anonymous):

x = no of coffee creamers left x + 47 = 100

OpenStudy (anonymous):

Thanks I know me and my husbane love your help haaaaaaaaaaaaaa @thejoker

OpenStudy (anonymous):

@Thejoker123

OpenStudy (anonymous):

Anytime. and your welcome